OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (lυἶცἶ0210):

Don't give away answers

OpenStudy (anonymous):

what?

OpenStudy (jdoe0001):

so... hmmm have you covered "solving by substitution" or "solving by elimination" yet?

OpenStudy (lulubj):

Solving by substitution will give you (1,3)

OpenStudy (anonymous):

Yes I read the lesson for that

OpenStudy (jdoe0001):

well.. .then you could just substitute say \(\bf 2x + y = 1\\ -4x + y = -1 \\ \quad \\ \quad \\ 2x+y=1\implies {\color{blue}{ y}}=1-2x\qquad then \\ \quad \\ -4x+{\color{blue}{ y}}=-1\implies ?\)

OpenStudy (lulubj):

Its really not that hard, if you need help on learning how to actually do it. I could help you

OpenStudy (anonymous):

O it just clicked Thanks everyone

OpenStudy (jdoe0001):

yw